# SIPDO: Closed-Loop Prompt Optimization via Synthetic Data Feedback

### v2

@@ Appendix A Optimzed Prompts for different tasks
our Task Requirements: Analyze a given premise (primary sentence) and determine whether it fully supports the truth of a hypothesis (subsequent sentence). Classify the relationship as either "Entailment" or "Non-Entailment" based on the logical and factual connections between the two. Provide the classification only as the final output. Problem Rule Application: Entailment: The premise explicitly confirms the hypothesis with clear, direct evidence. No additional information, assumptions, or interpretations are required to validate the hypothesis. Non-Entailment: The premise does not fully or explicitly confirm the hypothesis. If there is ambiguity, uncertainty, or missing logical links, label it as Non-Entailment. Judgment Criteria: (Strictly follow these rules) Language of Uncertainty: Words like "assumes," "believes," "thinks," "feels," "suspects" indicate subjectivity and should not be considered definitive proof. These terms suggest a possibility rather than an explicit factual connection. Specific vs. General Statements: A specific premise (e.g., mentioning a “full face mask”) does not necessarily contradict a general hypothesis (e.g., referencing a “mask” in general). However, if the premise is too general to confirm the specific claim, classify as Non-Entailment. Objective Reasoning: Only use the logical and factual ties within the given statements. Do not rely on external knowledge, assumptions, or interpretations unless directly supported by the premise. Decision Process: Determine whether the premise fully supports the hypothesis without needing extra inference → Entailment. If the premise only partially supports or fails to confirm the hypothesis → Non-Entailment. 0.893
@@
Table 9: Comparison of Optimized Prompts for Temporal task including CoT, APE, PromptAgent, and our method.
@@
CoT Your task is to determine the available time slot for an event, based on a schedule of occupied times. Follow these steps to ensure accuracy: **Steps to Identify Free Time Slots:** 1. **List Occupied Periods**: Organize all occupied time slots in chronological order. 2. **Find Gaps**: Identify gaps between the occupied periods where no activities are scheduled. 3. **Check Constraints**: Ensure that the free time slots fall within operational constraints (e.g., facility closing times). 4. **Select the Slot**: Choose the correct free time slot that satisfies all criteria. **Output Result Format:** - Present the selected free time slot in a clear format, such as "Xpm to Ypm" or "Xam to Yam". Use this step-by-step method to ensure that the identified time slot is accurate and does not overlap with any occupied periods. 0.989
@@
PromptAgent Analyze the timeline of events to determine possible time frames during which certain events could have occurred, even if they were not explicitly observed. Start by constructing a comprehensive timeline, clearly listing all observed and unobserved time slots. Identify gaps where the subject is unobserved, ensuring these gaps fit within any given constraints, such as opening and closing times. Emphasize the importance of constraints by verifying them after identifying potential gaps. Use a step-by-step reasoning approach to systematically evaluate all available information, and include a final review to check for potential errors or overlooked details before finalizing the answer. Define key terms like "unobserved" and "constraints" to ensure clarity in the task requirements. Provide examples to illustrate the reasoning process and expected output format, guiding the model in analyzing timelines and identifying possible time frames for unobserved events. Additionally, incorporate a checklist to ensure all steps are followed, and highlight common pitfalls to avoid during the analysis. Finally, include a summary of the reasoning process to reinforce understanding and ensure the model’s conclusions are well-supported. To further enhance the model’s performance, include additional examples that cover a wider range of scenarios and constraints, such as overlapping time slots or multiple constraints. Provide explicit guidance on handling complex constraints and ambiguous information. Incorporate interactive feedback mechanisms to help the model learn from mistakes and improve over time. Ensure the prompt is concise and focused, avoiding unnecessary repetition while maintaining clarity and comprehensiveness. Additionally, introduce a section for handling exceptions or unusual cases, offering strategies for dealing with incomplete or conflicting data. This will help the model adapt to a broader range of real-world scenarios and improve its robustness in timeline analysis tasks. 0.984
@@
our **Task Requirements:** Determine the possible time period during which an event could have occurred, based on a detailed schedule of occupied times. Your goal is to identify the correct time slot that fits all the provided criteria without any overlap. **Problem Rule Explanation:** 1. Analyze the schedule to identify all time slots during which the person is occupied. 2. Determine the available time slots by identifying gaps between these occupied periods. 3. Consider any additional constraints, such as closing times, that may limit the available time slots. **Problem Rule Application:** - List all the occupied time slots chronologically. - Identify gaps between these occupied slots where the person is free. - Ensure that the free time slots do not conflict with constraints like closing times. **Result Verification:** - Confirm that the identified time slot is completely free and adheres to any constraints. - Double-check against all occupied periods to ensure there is no overlap. - Avoid selecting time slots that are partially occupied or overlap with any scheduled activities. **Output Result Format:** - Present the correct time slot in a straightforward manner, using the format "Xpm to Ypm" or "Xam to Yam" as appropriate. - Ensure the output is clear and free of any extraneous symbols or text. **Common Mistakes to Avoid:** - Do not include time slots that extend beyond the closing time of the facility. - Avoid selecting time slots that overlap with any scheduled activities. - Ensure the selected time slot is entirely free and does not partially overlap with any occupied period. **General Rules and Analysis:** - Identify all occupied periods and list them chronologically. - Look for gaps between these periods where the person is not scheduled for any activity. - Verify that these gaps fall within any operational constraints, such as closing times. - Ensure the selected time slot is entirely free and does not overlap with any occupied periods. By following these guidelines, you can accurately determine the available time slot for the event in question. Avoid errors by ensuring that the selected time slot is entirely free and does not overlap with any occupied periods. 0.993
@@
Appendix B Optimized Prompts Without Difficulty Scaling in Synthetic Data
@@
Penguins You are provided with two tables containing data about penguins and giraffes. Your task is to focus solely on the giraffe data to answer a specific question regarding the tallest giraffe. **Penguin Data:** | Name | Age | Height (cm) | Weight (kg) | |———-|—–|————–|————-| | Louis | 7 | 50 | 11 | | Bernard | 5 | 80 | 13 | | Vincent | 9 | 60 | 11 | | Gwen | 8 | 70 | 15 | | James | 12 | 90 | 12 | **Giraffe Data:** | Name | Age | Height (cm) | Weight (kg) | |——–|—–|————–|————-| | Jody | 5 | 430 | 620 | | Gladys | 10 | 420 | 590 | | Marian | 2 | 310 | 410 | | Donna | 9 | 440 | 650 | **Task Requirements:** 1. Identify the tallest giraffe based on the height provided in the Giraffe Data table. 2. Provide the weight of the tallest giraffe in kilograms. **Problem Rule Explanation:** - Review the height values for each giraffe listed in the Giraffe Data table. - Compare these height values to determine which giraffe is the tallest. **Problem Rule Application:** - Examine the height values for the giraffes: - Jody: 430 cm - Gladys: 420 cm - Marian: 310 cm - Donna: 440 cm - Identify that Donna is the tallest giraffe at 440 cm. - Retrieve the corresponding weight of Donna, which is 650 kg. **Result Verification:** - Ensure that you have considered all entries in the Giraffe Data table. - Confirm that the weight you provide corresponds to the giraffe identified as the tallest. **Output Result Format:** - Provide your answer in the following format: - "Weight of the tallest giraffe: [Weight in kg]" **Example Output:** - "Weight of the tallest giraffe: 650" 0.732
Geometry " Given the following input: ""input"", you must provide ONLY the correct value for the ’target’. **Rules:** 1. Do NOT provide any explanations. 2. Do NOT provide any sentences, text, or words other than the ’target’ value. 3. The answer must be the exact value contained in the ""target"" and any unauthorized additions are prohibited." 0.681
Object Counting "**Task Requirements:** - Determine the total number of fruits by accurately identifying and counting each type listed in the question. **Problem Rule Explanation:** - The task involves listing and counting each fruit mentioned. - Each fruit should be counted as one unless a specific quantity is provided. **Problem Rule Application:** - Carefully read through the list to identify all items that are fruits. - Count each fruit once unless otherwise specified with a different quantity. - Avoid including any non-fruit items or miscounting due to misinterpretation of the list. **Result Verification:** - Re-examine the list to ensure all fruits have been correctly identified and counted. - Verify that the total count reflects only the fruits listed, with no errors in inclusion or exclusion. **Output Result Format:** - Provide the total number of fruits in both word and numeral forms, such as: [""ten"", ""10""]. - Ensure the output is clear and free from special symbols or formatting errors." 0.538
@@
Temporal **Task Requirements:** Determine the available time slots for an unscheduled activity within a given daily schedule, ensuring these slots do not conflict with scheduled events and comply with any facility operating hours. **Problem Rule Explanation:** 1. Review the entire schedule to identify all events and their specific time frames. 2. Identify gaps between these events or after the last scheduled event to find potential time slots for the unscheduled activity. 3. Consider any additional constraints, such as facility operating hours, to ensure the proposed time slot is feasible. **Problem Rule Application:** - List all scheduled events with their respective time frames. - Identify gaps between these events or available time after the last scheduled event. - Ensure that the identified time slots comply with any additional constraints, like facility operating hours. **Result Verification:** - Confirm that the identified time slots do not overlap with any scheduled events. - Verify that the time slots fall within the facility’s operating hours. **Output Result Format:** Present the time range in a clear and concise format, such as "Xpm to Ypm" or "Xam to Yam", ensuring clarity and precision. **Example Application:** Given the schedule: - Breakfast: 8am to 9am - Business meeting: 9am to 11am - Art gallery: 11am to 1pm - Lunch: 1pm to 2pm - Cinema: 3pm to 5pm - Dinner party: 6pm to 8pm - Gym closes at 10pm Determine the available time for the gym: - Identify the gaps: 5pm to 6pm and 8pm to 10pm. - Ensure these time slots do not overlap with scheduled events and are within the gym’s operating hours. - The correct answer is "5pm to 6pm" and "8pm to 10pm", as they fit within the gym’s operating hours and do not overlap with any scheduled events.
Appendix C Prompt Self-improvement
C.1 Error Analysis: Prompt Template
1. Misinterpretation of the Question: The model may have misinterpreted the question, focusing on the structure of the data rather than the specific request for the height of the tallest penguin. This could lead to confusion and an irrelevant answer.
@@ C.1 Error Analysis: Prompt Template
3. Irrelevant Output Generation: The answer "rectangle" does not relate to the context of the question. This suggests that the model may have generated a response based on unrelated patterns or associations rather than the specific data presented.
@@
C.2 Improvement Recommendation: Prompt Template
- Clarify the question to emphasize the need for numerical comparison.
- Provide explicit instructions to focus on extracting specific values from the data.
- Ensure the model is trained to recognize and process tabular data more effectively.
- Avoid ambiguity in the question to prevent misinterpretation of the request.
C.3 Improvement Recommendations: Revised Prompt
Here is a table of penguins with their respective attributes:
 name, age, height (cm), weight (kg)

### v3

@@ Appendix B Optimzed Prompts for different tasks
SIPDO Task Requirements: Analyze a given premise (primary sentence) and determine whether it fully supports the truth of a hypothesis (subsequent sentence). Classify the relationship as either ”Entailment” or ”Non-Entailment” based on the logical and factual connections between the two. Provide the classification only as the final output. Problem Rule Application: Entailment: The premise explicitly confirms the hypothesis with clear, direct evidence. No additional information, assumptions, or interpretations are required to validate the hypothesis. Non-Entailment: The premise does not fully or explicitly confirm the hypothesis. If there is ambiguity, uncertainty, or missing logical links, label it as Non-Entailment. Judgment Criteria: (Strictly follow these rules) Language of Uncertainty: Words like ”assumes,” ”believes,” ”thinks,” ”feels,” ”suspects” indicate subjectivity and should not be considered definitive proof. These terms suggest a possibility rather than an explicit factual connection. Specific vs. General Statements: A specific premise (e.g., mentioning a “full face mask”) does not necessarily contradict a general hypothesis (e.g., referencing a “mask” in general). However, if the premise is too general to confirm the specific claim, classify as Non-Entailment. Objective Reasoning: Only use the logical and factual ties within the given statements. Do not rely on external knowledge, assumptions, or interpretations unless directly supported by the premise. Decision Process: Determine whether the premise fully supports the hypothesis without needing extra inference → Entailment. If the premise only partially supports or fails to confirm the hypothesis → Non-Entailment. 0.893
@@
Table 14: Comparison of Optimized Prompts for Temporal task including CoT, APE, PromptAgent, and SIPDO.
@@
CoT Your task is to determine the available time slot for an event, based on a schedule of occupied times. Follow these steps to ensure accuracy: **Steps to Identify Free Time Slots:** 1. **List Occupied Periods**: Organize all occupied time slots in chronological order. 2. **Find Gaps**: Identify gaps between the occupied periods where no activities are scheduled. 3. **Check Constraints**: Ensure that the free time slots fall within operational constraints (e.g., facility closing times). 4. **Select the Slot**: Choose the correct free time slot that satisfies all criteria. **Output Result Format:** - Present the selected free time slot in a clear format, such as ”Xpm to Ypm” or ”Xam to Yam”. Use this step-by-step method to ensure that the identified time slot is accurate and does not overlap with any occupied periods. 0.989
@@
PromptAgent Analyze the timeline of events to determine possible time frames during which certain events could have occurred, even if they were not explicitly observed. Start by constructing a comprehensive timeline, clearly listing all observed and unobserved time slots. Identify gaps where the subject is unobserved, ensuring these gaps fit within any given constraints, such as opening and closing times. Emphasize the importance of constraints by verifying them after identifying potential gaps. Use a step-by-step reasoning approach to systematically evaluate all available information, and include a final review to check for potential errors or overlooked details before finalizing the answer. Define key terms like ”unobserved” and ”constraints” to ensure clarity in the task requirements. Provide examples to illustrate the reasoning process and expected output format, guiding the model in analyzing timelines and identifying possible time frames for unobserved events. Additionally, incorporate a checklist to ensure all steps are followed, and highlight common pitfalls to avoid during the analysis. Finally, include a summary of the reasoning process to reinforce understanding and ensure the model’s conclusions are well-supported. To further enhance the model’s performance, include additional examples that cover a wider range of scenarios and constraints, such as overlapping time slots or multiple constraints. Provide explicit guidance on handling complex constraints and ambiguous information. Incorporate interactive feedback mechanisms to help the model learn from mistakes and improve over time. Ensure the prompt is concise and focused, avoiding unnecessary repetition while maintaining clarity and comprehensiveness. Additionally, introduce a section for handling exceptions or unusual cases, offering strategies for dealing with incomplete or conflicting data. This will help the model adapt to a broader range of real-world scenarios and improve its robustness in timeline analysis tasks. 0.984
@@
SIPDO **Task Requirements:** Determine the possible time period during which an event could have occurred, based on a detailed schedule of occupied times. Your goal is to identify the correct time slot that fits all the provided criteria without any overlap. **Problem Rule Explanation:** 1. Analyze the schedule to identify all time slots during which the person is occupied. 2. Determine the available time slots by identifying gaps between these occupied periods. 3. Consider any additional constraints, such as closing times, that may limit the available time slots. **Problem Rule Application:** - List all the occupied time slots chronologically. - Identify gaps between these occupied slots where the person is free. - Ensure that the free time slots do not conflict with constraints like closing times. **Result Verification:** - Confirm that the identified time slot is completely free and adheres to any constraints. - Double-check against all occupied periods to ensure there is no overlap. - Avoid selecting time slots that are partially occupied or overlap with any scheduled activities. **Output Result Format:** - Present the correct time slot in a straightforward manner, using the format ”Xpm to Ypm” or ”Xam to Yam” as appropriate. - Ensure the output is clear and free of any extraneous symbols or text. **Common Mistakes to Avoid:** - Do not include time slots that extend beyond the closing time of the facility. - Avoid selecting time slots that overlap with any scheduled activities. - Ensure the selected time slot is entirely free and does not partially overlap with any occupied period. **General Rules and Analysis:** - Identify all occupied periods and list them chronologically. - Look for gaps between these periods where the person is not scheduled for any activity. - Verify that these gaps fall within any operational constraints, such as closing times. - Ensure the selected time slot is entirely free and does not overlap with any occupied periods. By following these guidelines, you can accurately determine the available time slot for the event in question. Avoid errors by ensuring that the selected time slot is entirely free and does not overlap with any occupied periods. 0.993
@@
Appendix C Optimized Prompts Without Difficulty Scaling in Synthetic Data
@@
Penguins You are provided with two tables containing data about penguins and giraffes. Your task is to focus solely on the giraffe data to answer a specific question regarding the tallest giraffe. **Penguin Data:** — Name — Age — Height (cm) — Weight (kg) — ————-——–—————–—————-— — Louis — 7 — 50 — 11 — — Bernard — 5 — 80 — 13 — — Vincent — 9 — 60 — 11 — — Gwen — 8 — 70 — 15 — — James — 12 — 90 — 12 — **Giraffe Data:** — Name — Age — Height (cm) — Weight (kg) — ———–——–—————–—————-— — Jody — 5 — 430 — 620 — — Gladys — 10 — 420 — 590 — — Marian — 2 — 310 — 410 — — Donna — 9 — 440 — 650 — **Task Requirements:** 1. Identify the tallest giraffe based on the height provided in the Giraffe Data table. 2. Provide the weight of the tallest giraffe in kilograms. **Problem Rule Explanation:** - Review the height values for each giraffe listed in the Giraffe Data table. - Compare these height values to determine which giraffe is the tallest. **Problem Rule Application:** - Examine the height values for the giraffes: - Jody: 430 cm - Gladys: 420 cm - Marian: 310 cm - Donna: 440 cm - Identify that Donna is the tallest giraffe at 440 cm. - Retrieve the corresponding weight of Donna, which is 650 kg. **Result Verification:** - Ensure that you have considered all entries in the Giraffe Data table. - Confirm that the weight you provide corresponds to the giraffe identified as the tallest. **Output Result Format:** - Provide your answer in the following format: - ”Weight of the tallest giraffe: [Weight in kg]” **Example Output:** - ”Weight of the tallest giraffe: 650” 0.732
Geometry ” Given the following input: ””input””, you must provide ONLY the correct value for the ’target’. **Rules:** 1. Do NOT provide any explanations. 2. Do NOT provide any sentences, text, or words other than the ’target’ value. 3. The answer must be the exact value contained in the ””target”” and any unauthorized additions are prohibited.” 0.681
Object Counting ”**Task Requirements:** - Determine the total number of fruits by accurately identifying and counting each type listed in the question. **Problem Rule Explanation:** - The task involves listing and counting each fruit mentioned. - Each fruit should be counted as one unless a specific quantity is provided. **Problem Rule Application:** - Carefully read through the list to identify all items that are fruits. - Count each fruit once unless otherwise specified with a different quantity. - Avoid including any non-fruit items or miscounting due to misinterpretation of the list. **Result Verification:** - Re-examine the list to ensure all fruits have been correctly identified and counted. - Verify that the total count reflects only the fruits listed, with no errors in inclusion or exclusion. **Output Result Format:** - Provide the total number of fruits in both word and numeral forms, such as: [””ten””, ””10””]. - Ensure the output is clear and free from special symbols or formatting errors.” 0.538
@@
Temporal **Task Requirements:** Determine the available time slots for an unscheduled activity within a given daily schedule, ensuring these slots do not conflict with scheduled events and comply with any facility operating hours. **Problem Rule Explanation:** 1. Review the entire schedule to identify all events and their specific time frames. 2. Identify gaps between these events or after the last scheduled event to find potential time slots for the unscheduled activity. 3. Consider any additional constraints, such as facility operating hours, to ensure the proposed time slot is feasible. **Problem Rule Application:** - List all scheduled events with their respective time frames. - Identify gaps between these events or available time after the last scheduled event. - Ensure that the identified time slots comply with any additional constraints, like facility operating hours. **Result Verification:** - Confirm that the identified time slots do not overlap with any scheduled events. - Verify that the time slots fall within the facility’s operating hours. **Output Result Format:** Present the time range in a clear and concise format, such as ”Xpm to Ypm” or ”Xam to Yam”, ensuring clarity and precision. **Example Application:** Given the schedule: - Breakfast: 8am to 9am - Business meeting: 9am to 11am - Art gallery: 11am to 1pm - Lunch: 1pm to 2pm - Cinema: 3pm to 5pm - Dinner party: 6pm to 8pm - Gym closes at 10pm Determine the available time for the gym: - Identify the gaps: 5pm to 6pm and 8pm to 10pm. - Ensure these time slots do not overlap with scheduled events and are within the gym’s operating hours. - The correct answer is ”5pm to 6pm” and ”8pm to 10pm”, as they fit within the gym’s operating hours and do not overlap with any scheduled events.
Appendix D Self-improvement Prompt templates
We provide an example of a self-improvement prompt template in the BIG-Bench Penguins In a Table task: Error analysis, Improvement recommendations, and prompt refinement.
Error Analysis Prompt Template You are an expert at analysing why LLM answer is wrong in Penguins in a table reasoning task. Your task is to give a concise and general description of the reasoning mistake from the current prompt that caused the mistake with the provided question, generated answer, ground truth, and current prompt below: Question :{synthetic question} Model generated answer: {LLM generated answer} Ground truth: {true answer} Current prompt: {current prompt} 1. Misinterpretation of the Question: The model may have misinterpreted the question, focusing on the structure of the data rather than the specific request for the height of the tallest penguin. This could lead to confusion and an irrelevant answer.
@@ Appendix D Self-improvement Prompt templates
3. Irrelevant Output Generation: The answer ”rectangle” does not relate to the context of the question. This suggests that the model may have generated a response based on unrelated patterns or associations rather than the specific data presented.
@@
Improvement Recommendation Prompt Template You are an expert in giving recommendations for optimizing current prompts. The goal is to output reasonable and decent suggestions on how to revise the current prompt to solve the encountered issue. Question :{synthetic question} Model generated answer: {LLM generated answer} Ground truth: {true answer} Current prompt: {current prompt} Generated error analysis: {error analysis from previous step} Some recommendation examples in Penguins In A Table task(But don’t be limited to these): - Clarify the question to emphasize the need for numerical comparison. - Provide explicit instructions to focus on extracting specific values from the data. - Ensure the model is trained to recognize and process tabular data more effectively. - Avoid ambiguity in the question to prevent misinterpretation of the request.
Prompt Refinement Template You are an expert in revising prompt engineering and refinement in tables related to reasoning tasks. Goal: Create a revised prompt that fixes the failure without overfitting – keep it generic. Question :{synthetic question} Model generated answer: {LLM generated answer} Ground truth: {true answer} Current prompt: {current prompt} Refinement recommendation: {recommendation from previous step} Tasks: - Give a better prompt, which can avoid all the problems that have occurred. - Summarize the logic based on the question and correct answer, and summarize the mistakes that should be avoided based on the question and generated wrong answer. - Do not modify the prompt based on the specific problem, but modify the prompt based on the cause of the error. The modified prompt should be able to give some regular hints and logical revisions. - You should specify the output format according to the correct answer. The output should not contain any special symbols. - You can summarize the characteristics of each option and reverse the answer based on the characteristics. - It can analyze the general rules of the problem, which helps the model understand how the problem is solved. - Prompts should be planned according to the following categories: task requirements, problem rule explanation, problem rule application, result verification, and output result format.
Appendix E Examples of Generated Data
Table 16: Examples of Generated Data for BIG-Bench tasks
Tasks Generated Data
Causal Judgment Question: In a small town, there is a bakery that makes the best pastries. Every morning, the bakery opens at 7:00 am, and a line of customers forms outside. The bakery owner has a rule that only one person can enter at a time to maintain order. One day, two customers, Alice and Bob, arrive at the same time. Alice follows the rule and waits outside, but Bob ignores the rule and enters the bakery while another customer is still inside. The bakery becomes overcrowded, and a shelf of pastries falls over, ruining the day’s batch. Did Bob cause the pastries to be ruined?
Answer: Yes
Geometry Question:This SVG path element <path d="M 50.00,30.00 L 66.18,35.09 L 72.45,50.00 L 66.18,64.91 L 50.00,70.00 L 33.82,64.91 L 27.55,50.00 L 33.82,35.09 L 50.00,30.00"/>
Answer: Polygon
Object Counting Question: I have two violins, a drum, a piano, a flute, and a trumpet. Additionally, I have a cat, a rabbit, a dog, a chicken, and a goat. How many musical instruments do I have?
Answer: 4
Epistemic Question: Premise: Olivia suspects that Ethan recognizes that a group of musicians gather in a park, tuning their instruments as the sun sets behind the city skyline. Hypothesis: Ethan recognizes that a group of musicians gather in a park, tuning their instruments as the sun sets behind the city skyline.
Answer: non-entailment
Temporal Question:Today, Alex attended several events. Between what times could he have gone to the gym? We know that: Alex had breakfast at 8am. He attended a meeting from 9am to 11am. He was seen at the art gallery from 11am to 1pm. He had lunch with friends from 1pm to 2pm. He was at the cinema from 2pm to 4pm. He visited his grandmother from 4pm to 6pm. The gym closes at 10pm. Between what times could Alex have gone to the gym? 12pm to 1pm, 9am to 10am, 6pm to 10pm, 4pm to 5pm
Answer: 6pm to 10pm
Penguins Question: Here is a table where the first line is a header and each subsequent line is a penguin: name, age, height (cm), weight (kg)